# Mr-Ben: A Comprehensive Meta-Reasoning Benchmark for Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have shown increasing capability in problem-solving and decision-making, largely based on the step-by-step chain-of-thought reasoning processes. However, it has been increasingly challenging to evaluate the reasoning capability of LLMs. Concretely, existing outcome-based benchmarks begin to saturate and become less sufficient to monitor the progress. To this end, we present a process-based benchmark Mr-Ben that demands a meta reasoning skill, where LMs are asked to locate and analyse potential errors in automatically generated reasoning steps. Mr-Ben is a comprehensive benchmark comprising 5,975 questions collected from human experts, covering various subjects such as physics, chemistry, logic, coding, and more. Through our designed metrics for assessing meta-reasoning on this benchmark, we identify interesting limitations and weaknesses of current LLMs (open-source and closed-source models). For example, open-source models are seemingly comparable to GPT-4 on outcome-based benchmarks, but they lag far behind on our benchmark, revealing the underlying reasoning capability gap between them. Our dataset and codes are available on https://randolph-zeng.github.io/Mr-Ben.github.io/.
@@ 1 Introduction
Reasoning, the cognitive process of using evidence, arguments, and logic to reach conclusions, is crucial for problem-solving, decision-making, and critical thinking [46, 15]. With the rapid advancement of Large Language Models (LLMs), there is an increasing interest in exploring their reasoning capabilities [20, 37]. Consequently, evaluating reasoning in LLMs reliably becomes paramount. Current evaluation methodologies primarily focus on the final result [12, 19, 16, 41], disregarding the intricacies of the reasoning process. While effective to some extent, such evaluation practices may conceal underlying issues like logical errors or unnecessary steps that compromise the accuracy and efficiency of reasoning [48].
Therefore, it is important to complement outcome-based evaluation with an intrinsic evaluation of the quality of the reasoning process. However, current benchmarks for evaluating LLMs’ reasoning capabilities have certain limitations in terms of their scope and size. For instance, PRM800K [25] categorizes each reasoning step as positive, negative, or neutral. Similarly, BIG-Bench Mistake [45] focuses on identifying errors in step-level answers. Other benchmarks like MR-GSM8K [50] and MR-Math [48] go a step further by providing the reason for the first negative step. However, these benchmarks have a restricted range of tasks, with MR-GSM8K and MR-Math concentrating solely on mathematics reasoning, while BIG-Bench Mistake mainly assesses logical reasoning. To ensure a comprehensive evaluation of reasoning abilities, it is crucial to identify reasoning errors and assess the LLMs’ capacity to elucidate them.
@@
Our analyses of various LLMs [32, 33, 5, 23, 30] reveal distinct limitations and previously unidentified weaknesses in their reasoning abilities. We find that while many LLMs can generate the correct answer to a question, they struggle to pinpoint errors in the reasoning process and correct them. This suggests that existing LLMs have yet to master reasoning, particularly the smaller models. However, we also find that techniques such as the use of high-quality synthetic data can significantly improve reasoning abilities, offering a potential pathway to enhance performance regardless of model size. Furthermore, we observe that different LLMs excel in different reasoning paradigms, challenging the assumption that domain-specific enhancements necessarily lead to broad cognitive improvements. We hope that Mr-Ben can guide the researchers to understand the abilities of their models from multiple dimensions and facilitate the development of AI reasoning frameworks. Our key contributions are summarized as follows:
We introduced Mr-Ben, which includes 6k questions across a wide range of subjects, from natural sciences to coding and logic, and employs a unique meta-reasoning paradigm.
We conduct an extensive analysis of various LLMs on Mr-Ben, revealing distinct limitations and previously unidentified weaknesses in their reasoning abilities.
We offer potential pathways for enhancing the reasoning abilities of LLMs and challenge the assumption that domain-specific enhancements necessarily lead to broad improvements,
@@ 3.1 Dataset Structure
To comprehensively evaluate the reasoning capabilities of LLMs, Mr-Ben employs a meta-reasoning paradigm. This paradigm casts LLMs in the role of a teacher, where they assess the reasoning process by evaluating its correctness, analyzing errors, and providing corrections. The construction pipeline is shown in Figure 6 in Appendix-C. As shown in Figure 1, each data point within Mr-Ben consists of three key elements: a question, a CoT answer, and an error analysis.
Question The questions in Mr-Ben are designed to cover a diverse range of reasoning types and difficulty levels, spanning from high school to professional levels. To ensure this breadth, we curated questions from various subjects, including natural sciences (mathematics, biology, physics), coding, and logic. Specifically, we sampled questions from mathematics, physics, biology, chemistry, and medicine from MMLU [18], which comprehensively assesses LLMs across academic and professional domains. For logic questions, we draw from LogiQA [27], which encompasses a broad spectrum of logical reasoning types, including categorical, conditional, disjunctive, and conjunctive reasoning. Finally, we select coding problems from MHPP [13], which focuses on function-level code generation requiring advanced algorithmic reasoning. Questions in MMLU and LogiQA require a single-choice answer, while MHPP requires a snippet of code as the answer.
CoT Answer We queried GPT-3.5-Turbo-0125 [32], Claude2 [5], and Mistral-Medium [22] (as of February 2024) using a prompt template (provided in Figure-7 in Appendix-C) designed to elicit step-by-step solutions [47]. For clarity, all LLMs were instructed to format their solutions with numbered steps, except for coding problems. To encourage diverse solutions, we set the temperature parameter to 1 during sampling. This empirical setting yielded satisfactory instruction following and desirable fine-grained reasoning errors, which annotators and evaluated models are expected to identify.
@@ 3.2 Annotation Process
Answer Correctness CoT answers that result in a final answer different from the ground truth are automatically flagged as incorrect. However, for cases where the final answer matches the ground truth, manual annotation is required. This is because there are instances where the reasoning process leading to the correct answer is flawed, as illustrated in the middle example of Figure 1. Therefore, annotators are tasked with meticulously examining the entire reasoning path to determine if the correct final answer is a direct result of the reasoning process.
Error Step This stage is applicable for solutions with either an unmatched final output or a matched final output underpinned by flawed reasoning. Following the prior effort [25], each step in the reasoning process is categorized as positive, neutral, or negative. Positive and neutral steps represent stages where the correct final output remains attainable. Conversely, negative steps indicate a divergence from the path leading to the correct solution. Annotators are required to identify the first step in the reasoning process where the conditions, assumptions, or calculations are incorrect, making the correct final result unreachable for the subsequent reasoning steps.
Error Reason and Correction Annotators are tasked with conducting an in-depth analysis of the reasoning that led to the identified error. As shown in Figure 1, annotators are required to provide the error reason and the corresponding correction to this reasoning step. This comprehensive approach ensures a thorough understanding and rectification of errors in the reasoning process.
@@ 3.4 Quality Control
Annotators Given the complexity of the questions, which span a range of subjects from high school to professional levels, we enlisted the services of an annotation company. This company meticulously recruited annotators, each holding a minimum of a bachelor’s degree. Before their trial labeling, annotators are thoroughly trained and are required to review the annotation guidelines. We’ve included the guidelines for all subjects in Appendix G for reference. The selection of annotators is based on their performance on a balanced, small hold-out set of problems for each subject. In addition to the annotators, a team of 14 quality controllers diligently monitors the quality of the annotation weekly. As a final layer of assurance, we have 4 meta controllers who scrutinize the quality of the work.
Quality Assurance Every problem in Mr-Ben undergoes a rigorous three-round quality assurance process to ensure its accuracy and clarity. Initially, each question is labeled by two different annotators. Any inconsistencies in the solution correctness or the first error step are identified and reviewed by a quality controller for arbitration. Following this, every annotated problem is subjected to a secondary review by annotators who were not involved in the initial labeling. This is to ensure that the annotations for different solutions to the same problem are consistent and coherent. In the final phase of the review, 10% of the problems are randomly sampled and reviewed by the meta controllers. Throughout the entire evaluation process, all annotated fields are meticulously examined in multiple rounds for their accuracy and clarity. Any incorrect annotations or those with disagreements are progressively filtered out and rectified, ensuring a high-quality dataset. This rigorous process allows us to maintain a high level of annotation quality.
Dataset Artifacts & Biases Table 1 reveals a relatively balanced distribution of correct and incorrect solutions. However, an exception was observed in mathematical subjects, where the distribution tends to skew towards incorrect solutions. This skew could suggest an inherent complexity or ambiguity in mathematical problem statements. Our analysis of the first error step across all subjects indicated that errors predominantly occur in the initial stages () of problem-solving and are distributed relatively uniformly. This pattern was consistent across most subjects, with no significant skew towards later steps. More detailed discussions of biases are provided in the Appendix B.
@@ 4 Evaluation
For each question-solution pair annotated, the evaluated model are supposed to decide the correctness of the solution and report the first-error-step and error-reason if any. The solution-correctness and first-error-step is scored automatically based on the manual annotation result. Only when the evaluated model correctly identified the incorrect solution and first-error-step will its error-reason be further examined manually or automatically by models. Therefore in order to provide a unified and normalized score to reflect the overall competence of the evaluated model, we follow the work of [50] and apply metric named MR-Score, which consist of three sub-metrics.
The first one is the Matthews Correlation Coefficient (a.k.a MCC, 29) for the binary classification of solution-correctness.
@@
For the weights and , they are chosen based on our evaluation results to maximize the differentiation between different models. It is important to note that the Matthews Correlation Coefficient (MCC) and the accuracy of locating the first error step can be directly calculated by comparing the responses of the evaluated model with the ground truth annotations. However, assessing the accuracy of the error reason explained by the evaluated model presents more complexity. While consulting domain experts for annotations is a feasible approach, we instead utilized GPT-4 as a proxy to examine the error reasons, as detailed in Figure-11 in Appendix-C. We operate under the assumption that while our benchmark presents a significant challenge for GPT-4 in evaluating complete solution correctness—identifying the first error step and explaining the error reason—it is comparatively easier for GPT-4 to assess whether the provided error reasons align with the ground truth. Specifically, in a hold-out set of sampled error reasons, there was a 92% agreement rate between the manual annotations by the authors and those generated by GPT-4.
@@ 5.1 Experiment Setup
To evaluate the performance of different models on our new benchmark, we selected a diverse array of models based on size and source accessibility 111 Note: All models used in our experiments are instruction-finetuned versions, although this is not indicated in their abbreviated names. This included smaller models like Gemma-2B[44], Phi-3[1], Qwen1.5-1.8B [7], as well as larger counterparts such as Llama3-70B [30], Deepseek-67B[8], and Qwen1.5-72B[7]. We also compared open-source models (e.g. models from the Llama3 and Qwen1.5/Qwen2 series) against closed-source models from the GPT [33], Claude [6], Mistral [22], GLM [3], Yi [26], Moonshot [2], Doubao [9] families. Additionally, models from the Deepseek-Coder [8] series were included to assess the impact of coding-focused pretraining on reasoning performance.
Given the complexity of our benchmark, even larger open-source models like Llama3-70B-Instruct struggle to produce accurate evaluation results without the use of prompting methods, often achieving MR-Scores near zero. Consequently, we employed a step-wise chain-of-thought prompting technique similar to those described in [50, 45]. This approach guides models in systematically reasoning through solution traces before making final decisions, as detailed in Appendix-C.
Considering the complexity of the task, which includes question comprehension, reasoning through the provided solutions and adhering to format constraints, few-shot demonstration setups are also explored to investigate if models can benefit from In-Context Learning (ICL) examples. Due to the context token limits, we report zero and one-shot results in the main result table (Table-2). The performance of additional few-shot configurations on a selection of models with various capabilities is further discussed in Section 6.1.
@@ 5.2 Experiment Results
The Mr-Ben benchmark presents a significant shift in challenge for state-of-the-art large language models, transitioning from question-answering to the nuanced role of question-solution scoring. This section details our findings, emphasizing variations in model performances and their implications.
@@
Overall Model Performance: Among the evaluated models, GPT-4-turbo consistently achieves the highest MR-Scores across all subjects, significantly outperforming most competitors from both open and closed-source communities. Notably, the open sourced Qwen2-72B and Deepseek-V2-236B models are performing exceptionally well, surpassing every other open sourced models including Llama3 by large margin. Their scores are even comparable or greater than some of the most capable models from the commercial companies, such as Mistral, Yi,and Moonshot AI. In the small language model category, the performance of Phi3-3.8B exceeds many from the mid-size models, including Deepseek-Coder-33B, whose size is around tenfold larger (see Table-2 and Figure-3).
@@
Performance by Model Size: Results from Table-2 reveals a general trend where larger models tend to perform better, highlighting the correlation between model size and the efficacy in complex reasoning tasks. However, this relationship is not strictly linear, as demonstrated by models like Phi3-3.8B, which excel despite their smaller size. This suggests that while larger model sizes generally yield superior performances, efficient data synthesis techniques like knowledge distillation can also significantly influence outcomes.
@@
Categorization by Reasoning Paradigm: Our categorization into four reasoning types—knowledge, arithmetic, algorithmic, and logic—illustrates the unique challenges each model faces within these paradigms (Figure-3). Logic reasoning emerges as the most formidable due to the intricate logical operations required by questions from the LogiQA dataset. In stark contrast, GPT-4-turbo and Llama3-70B demonstrate exceptional prowess in algorithmic reasoning, where their capabilities markedly surpass other models. Notably, models excel in different reasoning paradigms, reflecting their varied strengths and training backgrounds. For instance, despite Deepseek-Coder’s specialized pre-training focused on coding tasks, it does not necessarily confer superior abilities in algorithmic reasoning, underscoring that targeted pretraining does not guarantee enhanced performance across all reasoning types. Comparing the performance of Deepseek-Coder with that of the Phi-3 model, which excels despite much smaller size, highlights the potential significance of high-quality synthetic data in achieving broad-based reasoning capabilities.
@@
Sensitivity to Task Difficulty and Solution Length: An examination across educational levels shows some models perform slightly better on college level questions and some slightly better on high school level difficulties, indicating a low sensitivity to the difficulty levels of the questions. Additionally, our analysis find a minor negative correlation between the length of solution steps and MR-Scores, as detailed in Figure-5 and Figure-5.
@@
Summary: The Mr-Ben benchmark effectively differentiates model capabilities, often obscured in simpler settings. It not only identifies top performers but also underscores the influence of model size on outcomes, while demonstrating that techniques like knowledge distillation, as seen with the Phi-3 model, can notably enhance smaller models’ performance, challenging the dominance of larger models. The analysis further reveals that specialized training, such as in coding, does not guarantee superior algorithmic reasoning. This suggests the potential need for more balanced data approaches or improved data synthesis methods.

### v2

# MR-Ben: A Meta-Reasoning Benchmark for Evaluating System-2 Thinking in LLMs

@@ Abstract
Large language models (LLMs) have shown increasing capability in problem-solving and decision-making, largely based on the step-by-step chain-of-thought reasoning processes. However, evaluating these reasoning abilities has become increasingly challenging. Existing outcome-based benchmarks are beginning to saturate, becoming less effective in tracking meaningful progress. To address this, we present a process-based benchmark MR-Ben that demands a meta-reasoning skill, where LMs are asked to locate and analyse potential errors in automatically generated reasoning steps. Our meta-reasoning paradigm is especially suited for system-2 slow thinking, mirroring the human cognitive process of carefully examining assumptions, conditions, calculations, and logic to identify mistakes. MR-Ben comprises 5,975 questions curated by human experts across a wide range of subjects, including physics, chemistry, logic, coding, and more. Through our designed metrics for assessing meta-reasoning on this benchmark, we identify interesting limitations and weaknesses of current LLMs (open-source and closed-source models). For example, with models like the o1 series from OpenAI demonstrating strong performance by effectively scrutinizing the solution space, many other state-of-the-art models fall significantly behind on MR-Ben, exposing potential shortcomings in their training strategies and inference methodologies111Our dataset and codes are available on https://randolph-zeng.github.io/Mr-Ben.github.io..
@@ 1 Introduction
Reasoning, the cognitive process of using evidence, arguments, and logic to reach conclusions, is crucial for problem-solving, decision-making, and critical thinking [65, 19]. With the rapid advancement of Large Language Models (LLMs), there is an increasing interest in exploring their reasoning capabilities [30, 57]. Consequently, evaluating reasoning in LLMs reliably becomes paramount. Current evaluation methodologies primarily focus on the final result [16, 28, 22, 60], disregarding the intricacies of the reasoning process. While effective to some extent, such evaluation practices may conceal underlying issues like logical errors or unnecessary steps that compromise the accuracy and efficiency of reasoning [68, 41].
Therefore, it is important to complement outcome-based evaluation with an intrinsic evaluation of the quality of the reasoning process. However, current benchmarks for evaluating LLMs’ reasoning capabilities have certain limitations in terms of their scope and size. For instance, PRM800K [38] categorizes each reasoning step as positive, negative, or neutral. Similarly, BIG-Bench Mistake [64] focuses on identifying errors in step-level answers. We follow the same meta-reasoning paradigm as MR-GSM8K [77] and MR-Math [68], which go a step further by providing the error reason for the first negative step in the reasoning chain. However, these benchmarks are limited to a narrower task scope—MR-GSM8K and MR-Math focus solely on mathematical reasoning, while BIG-Bench Mistake mainly assesses logical reasoning. To ensure a comprehensive evaluation of reasoning abilities, it is crucial to identify reasoning errors and assess the LLMs’ capacity to elucidate them across wider domains.
@@
Our analysis of various LLMs [50, 51, 5, 33, 47] uncovers distinct limitations and previously unidentified weaknesses in their reasoning abilities. While many LLMs are capable of generating correct answers, they often struggle to identify errors within their reasoning processes and explain the underlying rationale. To excel under our meta-reasoning paradigm, models must meticulously scrutinize assumptions, conditions, calculations, and logical steps, even inferring step outcomes counterfactually. These requirements align with the characteristics of “System-2” slow thinking [35, 9], which we believe remains underdeveloped in most of the state-of-the-art models we evaluated.
We suspect that a key reason for this gap lies in current fine-tuning paradigms, which prioritize correct solutions and limit effective exploration of the broader solution space. Echoing this hypothesis, we observed that models like o1-preview [52], which reportedly incorporate effective search and disambiguation techniques across trajectories in the solution space, outperform other models by a large margin. Moreover, we found that leveraging high-quality and diverse synthetic data [1] significantly mitigates this issue, offering a promising path to enhance performance regardless of model size. Additionally, our results indicate that different LLMs excel in distinct reasoning paradigms, challenging the notion that domain-specific enhancements necessarily yield broad cognitive improvements. We hope that MR-Ben will guide researchers in comprehensively evaluating their models’ capabilities and foster the development of more robust AI reasoning frameworks.
Our key contributions are summarized as follows:
We introduced MR-Ben, which includes around 6k questions across a wide range of subjects, from natural sciences to coding and logic, and employs a unique meta-reasoning paradigm.
We conduct an extensive analysis of various LLMs on MR-Ben, revealing various limitations and previously unidentified weaknesses in their reasoning abilities.
We offer potential pathways for enhancing the reasoning abilities of LLMs and challenge the assumption that domain-specific enhancements necessarily lead to broad improvements.
@@ 3.1 Dataset Structure
To comprehensively evaluate the reasoning capabilities of LLMs, MR-Ben employs a meta-reasoning paradigm. This paradigm casts LLMs in the role of a teacher, where they assess the reasoning process by evaluating its correctness, analyzing errors, and providing corrections. As shown in Figure 1, each data point within MR-Ben consists of three key elements: a question, a CoT answer, and an error analysis. The construction pipeline is shown in Figure 6 in Appendix-D.
Question
The questions in MR-Ben are designed to cover a diverse range of reasoning types and difficulty levels, spanning from high school to professional levels. To ensure this breadth, we curated questions from various subjects, including natural sciences (mathematics, biology, physics), coding, and logic. Specifically, we sampled questions from mathematics, physics, biology, chemistry, and medicine from MMLU [27], which comprehensively assesses LLMs across academic and professional domains. For logic questions, we draw from LogiQA [40], which encompasses a broad spectrum of logical reasoning types, including categorical, conditional, disjunctive, and conjunctive reasoning. Finally, we select coding problems from MHPP [17], which focuses on function-level code generation requiring advanced algorithmic reasoning. Questions in MMLU and LogiQA require a single-choice answer, while MHPP requires a snippet of code as the answer.
CoT Answer
We queried GPT-3.5-Turbo-0125 [50], Claude2 [5], and Mistral-Medium [32] (as of February 2024) using a prompt template (provided in Figure-7 in Appendix-D) designed to elicit step-by-step solutions [66]. For clarity, all LLMs were instructed to format their solutions with numbered steps, except for coding problems. To encourage diverse solutions, we set the temperature parameter to 1 during sampling. This empirical setting yielded satisfactory instruction following and desirable fine-grained reasoning errors, which annotators and evaluated models are expected to identify.
@@ 3.2 Annotation Process
Answer Correctness
CoT answers that result in a final answer different from the ground truth are automatically flagged as incorrect. However, for cases where the final answer matches the ground truth, manual annotation is required. This is because there are instances where the reasoning process leading to the correct answer is flawed, as illustrated in the middle example of Figure 1. Therefore, annotators are tasked with meticulously examining the entire reasoning path to determine if the correct final answer is a direct result of the reasoning process.
Error Step
This stage is applicable for solutions with either an unmatched final output or a matched final output underpinned by flawed reasoning. Following the prior effort [38], each step in the reasoning process is categorized as positive, neutral, or negative. Positive and neutral steps represent stages where the correct final output remains attainable. Conversely, negative steps indicate a divergence from the path leading to the correct solution. Annotators are required to identify the first step in the reasoning process where the conditions, assumptions, or calculations are incorrect, making the correct final result unreachable for the subsequent reasoning steps.
Error Reason and Correction
Annotators are tasked with conducting an in-depth analysis of the reasoning that led to the identified error. As shown in Figure 1, annotators are required to provide the error reason and the corresponding correction to this reasoning step. This comprehensive approach ensures a thorough understanding and rectification of errors in the reasoning process.
@@ 3.4 Quality Control
Annotators
Given the complexity of the questions, which span a range of subjects from high school to professional levels, we enlisted the services of an annotation company. This company meticulously recruited annotators, each holding a minimum of a bachelor’s degree. Before their trial labeling, annotators are thoroughly trained and are required to review the annotation guidelines. We’ve included the guidelines for all subjects in Appendix H for reference. The selection of annotators is based on their performance on a balanced, small hold-out set of problems for each subject. In addition to the annotators, a team of 14 quality controllers diligently monitors the quality of the annotation weekly. As a final layer of assurance, we have 4 meta controllers who scrutinize the quality of the work.
Quality Assurance
Every problem in MR-Ben undergoes a rigorous three-round quality assurance process to ensure its accuracy and clarity. Initially, each question is labeled by two different annotators. Any inconsistencies in the solution correctness or the first error step are identified and reviewed by a quality controller for arbitration. Following this, every annotated problem is subjected to a secondary review by annotators who were not involved in the initial labeling. This is to ensure that the annotations for different solutions to the same problem are consistent and coherent. In the final phase of the review, 10% of the problems are randomly sampled and reviewed by the meta controllers. Throughout the entire evaluation process, all annotated fields are meticulously examined in multiple rounds for their accuracy and clarity. Any incorrect annotations or those with disagreements are progressively filtered out and rectified, ensuring a high-quality dataset. This rigorous process allows us to maintain a high level of annotation quality.
Dataset Artifacts & Biases
Table 1 reveals a relatively balanced distribution of correct and incorrect solutions. However, an exception was observed in mathematical subjects, where the distribution tends to skew towards incorrect solutions. This skew could suggest an inherent complexity or ambiguity in mathematical problem statements. Our analysis of the first error step across all subjects indicated that errors predominantly occur in the initial stages () of problem-solving and are distributed relatively uniformly. This pattern was consistent across most subjects, with no significant skew towards later steps. More detailed discussions of biases are provided in the Appendix C.
@@ 4 Evaluation
For each question-solution pair annotated, the evaluated model are supposed to decide the correctness of the solution and report the first-error-step and error-reason if any. The solution-correctness and first-error-step is scored automatically based on the manual annotation result. Only when the evaluated model correctly identified the incorrect solution and first-error-step will its error-reason be further examined manually or automatically by models. Therefore in order to provide a unified and normalized score to reflect the overall competence of the evaluated model, we follow the work of [77] and apply a metric named MR-Score, which consist of three sub-metrics.
The first one is the Matthews Correlation Coefficient (a.k.a MCC, 46) for the binary classification of solution-correctness.
@@
For the weights and , they are chosen based on our evaluation results to maximize the differentiation between different models. It is important to note that the Matthews Correlation Coefficient (MCC) and the accuracy of locating the first error step can be directly calculated by comparing the responses of the evaluated model with the ground truth annotations. However, assessing the accuracy of the error reason explained by the evaluated model presents more complexity. While consulting domain experts for annotations is a feasible approach, we instead utilized GPT-4-Turbo as a proxy to examine the error reasons, as detailed in Figure-11 in Appendix-D.
We operate under the assumption that while our benchmark presents a significant challenge for GPT-4 in evaluating complete solution correctness—identifying the first error step and explaining the error reason—it is comparatively easier for GPT-4 to assess whether the provided error reasons align with the ground truth. Specifically, in a hold-out set of sampled error reasons, there was a 92% agreement rate between the manual annotations by the authors and those generated by GPT-4. For more detailed evaluations on the robustness of MR-Score and its design thinking, please refer to our discussion in Appendix-B.
@@ 5.1 Experiment Setup
To evaluate the performance of different models on our new benchmark, we selected a diverse array of models based on size and source accessibility 222 Note: All models used in our experiments are instruction-finetuned versions, although this is not indicated in their abbreviated names. This included smaller models like Gemma-2B[63], Phi-3[1], Qwen1.5-1.8B [7], as well as larger counterparts such as Llama3-70B [47], Deepseek-67B[10], and Qwen1.5-72B[7]. We also compared open-source models (e.g. models from the Llama3 and Qwen1.5/Qwen2 series) against closed-source models from the GPT [51], Claude [6], Mistral [32], GLM [3], Yi [39], Moonshot [2], Doubao [12] families. Additionally, models from the Deepseek-Coder [10] series were included to assess the impact of coding-focused pretraining on reasoning performance.
Given the complexity of our benchmark, even larger open-source models like Llama3-70B-Instruct struggle to produce accurate evaluation results without the use of prompting methods, often achieving MR-Scores near zero. Consequently, we employed a step-wise chain-of-thought prompting technique similar to those described in [77, 64]. This approach guides models in systematically reasoning through solution traces before making final decisions, as detailed in Appendix-D.
Considering the complexity of the task, which includes question comprehension, reasoning through the provided solutions, and adhering to format constraints, few-shot demonstration setups are also explored to investigate if models can benefit from In-Context Learning (ICL) examples. Due to the context token limits, we report zero and one-shot results in the main result table (Table 2)333For the breakdown performances of models in the sub-tasks, please refer to Table 7. The performance of additional few-shot configurations on a selection of models with various capabilities is further discussed in Section 6.1.
@@ 5.2 Experiment Results
The MR-Ben benchmark presents a significant shift in the challenge for state-of-the-art large language models, transitioning from question-answering to the nuanced role of question-solution scoring. This section details our findings, emphasizing variations in model performances and their implications.
Overall Performance
@@
Among the evaluated models, o1-preview consistently achieves the highest MR-Scores across all subjects, significantly outperforming most competitors from both open and closed-source communities. Notably, the open-sourced Qwen2-72B and Deepseek-V2-236B models are performing exceptionally well, surpassing every other open-sourced model including Llama3 by a large margin. Their scores are even comparable to or greater than some of the most capable models from commercial companies, such as Mistral, Yi, and Moonshot AI. In the small language model category, the performance of Phi3-3.8B exceeds many of the mid-size models, including Deepseek-Coder-33B, whose size is around tenfold larger.
Performance across Model Size and Reasoning Paradigm
@@
Table 2 reveals a general trend where larger models tend to perform better, highlighting the correlation between model size and the efficacy in complex reasoning tasks. However, this relationship is not strictly linear, as demonstrated by models like Phi3-3.8B, which excel despite their smaller size. Since MR-Ben challenges the language models to reason about the reasoning in the solution space among a diverse range of domains, models like Phi-3 that are trained with effective data synthesis techniques and broader coverage of the solution space, intuitively achieve higher MR-Score. This suggests that while larger model sizes generally yield superior performances, techniques like knowledge distillation can also significantly boost reasoning performance. Similarly, although the size of the o1 model series remains undisclosed, these models reportedly employ mechanisms that scale computation efficiently through effective exploration, frequent retrospection, and meticulous reflection within the solution space. These characteristics align closely with the principles of “system-2” thinking, which emphasizes deliberate, reflective problem-solving. As a result, the o1 models demonstrate a more effective reasoning process, achieving significantly higher MR-Scores than other models by a large margin.
Performance across Reasoning Types
@@
Our categorization into four reasoning types—knowledge, arithmetic, algorithmic, and logic—illustrates the unique challenges each model faces within these paradigms (Figure 3). Logic reasoning emerges as the most formidable due to the intricate logical operations required by questions from the LogiQA dataset. In stark contrast, o1-Preview and GPT-4-turbo demonstrate exceptional prowess in algorithmic reasoning, where their capabilities markedly surpass other models. Notably, models excel in different reasoning paradigms, reflecting their varied strengths and training backgrounds. For instance, despite Deepseek-Coder’s specialized pre-training focused on coding tasks, it does not necessarily confer superior abilities in algorithmic reasoning, underscoring that targeted pretraining does not guarantee enhanced performance across all reasoning types. Comparing the performance of the Deepseek-Coder with that of the Phi-3 model, which excels despite its much smaller size, highlights the potential significance of high-quality synthetic data in achieving broad-based reasoning capabilities.
Sensitivity to Task Difficulty and Solution Length
@@
An examination across educational levels shows most models perform better at high school-level questions than college-level ones, indicating an intuitive level of sensitivity to the difficulty levels of the questions. Additionally, our analysis finds a minor negative correlation between the length of solution steps and MR-Scores, as detailed in Figure 5 and Figure 5.
Summary:
@@
MR-Ben effectively differentiates model capabilities, often obscured in simpler settings. It not only identifies top performers but also underscores the influence of model size on outcomes, while demonstrating that techniques like knowledge distillation and test-time compute scaling, as seen with the Phi-3 and o1 models, can notably enhance smaller models’ performance, challenging the dominance of larger models. The analysis further reveals that specialized training, such as in coding, does not guarantee superior algorithmic reasoning. This suggests the potential need for more balanced data approaches or improved data synthesis methods.